OpenStudy (jaydelv):

There are 40 volunteers for the research study on the Power Pill. Each subgroup of the study will contain 10 participants. Determine how many ways these participants can be selected and explain your method.

OpenStudy (anonymous):

(40*39*38......*30)*4

Vocaloid (vocaloid):

use the combination formula nCr = n!/(r!*(n-r)!) where n = 40 and r = 10

OpenStudy (jaydelv):

Can you explain the ! to me?

Vocaloid (vocaloid):

! means factorial examples: 5! = 5*4*3*2*1 3! = 3*2*1 and so on...

OpenStudy (anonymous):

How are you asking such question and you do not know what does factorial mean. Why do you go to school man stay at home

OpenStudy (jaydelv):

I'm taking a virtual course in a foreign country right now and I'm on a deadline so I'm rushing through this. I don't need your slick remarks because clearly you are seeing me trying to learn how to do this.

OpenStudy (anonymous):

How old are you?

OpenStudy (jaydelv):

None of your concern.

OpenStudy (jaydelv):

@Vocaloid I can't find the answer to the solution

OpenStudy (jaydelv):

Equation *

OpenStudy (anonymous):

Sorry for bothering you but i cant endure fool in maths its not in my control sorry bro

OpenStudy (anonymous):

Fools*

OpenStudy (jaydelv):

Grow up